OpenStudy (anonymous):

How to write 3/a-2 + 1/a^2-4 as a single fraction?

Parth (parthkohli):

\[\begin{align}= \rm{3 \over a - 2} + {1 \over a^2 -4} \\ \\ \rm= {3 \over a -2}+ {1 \over (a + 2)(a - 2)} \\ \\\rm = {3(a + 2) \over (a - 2)(a + 2)} + {1 \over (a + 2)(a - 2)} \\ \\\rm = {3a + 6 + 1 \over (a + 2)(a - 2)} \\ \\=\rm {3a + 7 \over a^2 - 4}\end{align} \]

Parth (parthkohli):

1) Get a common denominator. 2) Operate.

Parth (parthkohli):

For a little quiz, @mel2, how do you write this?\[{3 \over \rm 2a} + \rm{9219323838 \over a}\]

Parth (parthkohli):

I'm afraid you need a little technological help at this moment—unless you are a supercomputer!